OpenStudy (anonymous):

write answer in a/b form where a and b are integers. 27^-2/3

OpenStudy (mertsj):

\[27^{\frac{-2}{3}}= \frac{1}{27^{\frac{2}{3}}}=\frac{1}{(\sqrt[3]{27})^2}=\frac{1}{3^2}=\frac{1}{9}\]

OpenStudy (anonymous):

why did you have to make it into fraction like that? was it just to make it postive